# Degree of Kripke-incompleteness of Tense Logics

### v1

@@ 0.4.1 Degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K}_{t})$
[…]
@@
Towards a contradiction, suppose there exists a logic $L$ splits $\mathsf{NExt}(L^{*})$. Since $L\supseteq L^{*}$, $L\nsubseteq\mathsf{Log}(\bullet)$. By Theorem 0.2.31, $L$ does not split $\mathsf{NExt}(\mathsf{K}_{t})$. By Fact 0.2.30, $\bigcap_{i\in I}L_{i}\subseteq L$ for a family of logics $\mathcal{K}={\{L_{i}\in\mathsf{NExt}(\mathsf{K}_{t}):L_{i}\nsubseteq L,i\in I\}}$. Let $\mathcal{K}^{\prime}={\{L_{i}\oplus(\Diamond\top\vee\blacklozenge\top):i\in I\}}$. Take any $\varphi\not\in L$. Since $L^{*}\subseteq L$, $(\Diamond\top\vee\blacklozenge\top)\to\varphi\not\in L$. Since $\bigcap_{i\in I}L_{i}\subseteq L$, $(\Diamond\top\vee\blacklozenge\top)\to\varphi\not\in L_{i}$ for some $i\in I$. Then there exists $\mathbb{F}\in\mathsf{GF}(L_{i})$ such that $\mathbb{F}\not\models(\Diamond\top\vee\blacklozenge\top)\to\varphi$. By Proposition 0.4.1, it is not hard to see that $\mathbb{F}\models(\Diamond\top\vee\blacklozenge\top)$ and $\mathbb{F}\not\models\varphi$. Thus $\varphi\not\in L_{i}\oplus(\Diamond\top\vee\blacklozenge\top)$. Since $\varphi$ is arbitrarily chosen, we see that $\bigcap\mathcal{K}^{\prime}\subseteq L$. Note that $L$ splits $\mathsf{NExt}(L^{*})$ and $L\subseteq L_{i}\oplus L\nsubseteq L_{1}$ for any $i\in I$, we see that $\bigcap_{i\in I}L_{i}\oplus L\nsubseteq L_{1}$, which leads to a contradiction. ∎
@@
For (1), we prove by induction on $n$. Let $n=0$. Note that $\breve{R_{I}}[0]=\varnothing$, $k\in R_{I}^{k}[0]$ and $\breve{R_{I}}^{k+1}[k]=\varnothing$, we have $\mathbb{F}_{I},0\models\gamma_{0}$. Suppose $\mathbb{F}_{I},x\models\gamma_{0}$. Then $\mathbb{F}_{I},x\models\Diamond^{k}\blacksquare^{k+1}\bot$, which entails that there exists a strict chain ${\langle x_{i}:i\leq k\rangle}$ with $x=x_{0}$. Since $|\mathfrak{F}_{L}|<k$, we see $x\not\in X_{L}$. Since $\mathbb{F}_{I},x\models\blacksquare\bot\wedge\Diamond\blacksquare^{2}\bot$, we have $x\not\in\mathbb{Z}^{+}\cup{\{i^{*}:i\in I\}}$. Thus $x=0$. Let $n>0$. By induction hypothesis, for all $y\in X_{I}$, $\mathbb{F}_{I},y\models\gamma_{n-1}$ if and only if $y=n-1$. Since $n-1\in\breve{R_{I}}[n]\setminus\breve{R_{I}}^{2}[n]$, we have $\mathbb{F}_{I},n\models\gamma_{n}$. Suppose $\mathbb{F}_{I},x\models\gamma_{n}$. By induction hypothesis, $x\in R_{I}[n-1]\setminus R_{I}^{2}[n-1]$, which entails $x=n$.
@@
Suppose there exists $\mathfrak{G}=(Y,S)\in\mathsf{Fr}(\mathsf{Log}(\mathbb{F}_{I}))$. By an easy induction, we see that for all $V\in A_{I}$, $V\cap\omega$ and $V\cap{\{i^{*}:i\in I\cup{\{0\}}\}}$ are either finite or co-finite. Thus $\mathbb{F}_{I},1\models\Box(\Box p\to p)\to\Box p$. By Lemma 0.4.10(1), $\mathbb{F}_{I}\models\gamma_{i}\to\Diamond\gamma_{i+1}$ for all $i\in\omega$. Thus we have
$\mathfrak{G}\models\gamma_{1}\to(\Box(\Box p\to p)\to\Box p)$ and $\mathfrak{G}\models{\{\gamma_{i}\to\Diamond\gamma_{i+1}:i\in\omega\}}$.
Note that $\mathbb{F}_{I}\models\Delta^{k}\gamma_{1}$ and $\gamma_{1}$ is variable-free, there exists $y\in Y$ such that $\mathfrak{G},y\models\gamma_{1}$. Since $\mathfrak{G}\models{\{\gamma_{i}\to\Diamond\gamma_{i+1}:i\in\omega\}}$, we see that there exists an infinite strictly ascending $S$-chain ${\langle u_{i}:i\in\mathbb{Z}^{+}\rangle}$ such that $y=u_{1}$ and $\mathfrak{G},u_{i}\models\gamma_{i}$ for all $i\in\mathbb{Z}^{+}$. Let $U$ be a valuation in $\mathfrak{G}$ such that $U(p)={\{u_{2i}:i\in\mathbb{Z}^{+}\}}$. Then we see that $\mathfrak{G},U,y\not\models\Box(\Box p\to p)\to\Box p$. Hence $\mathfrak{G}\not\models\gamma_{1}\to(\Box(\Box p\to p)\to\Box p)$, which contradicts $\mathfrak{G}\in\mathsf{Fr}(\mathsf{Log}(\mathbb{F}_{I}))$. ∎
@@
For all $I,J\in\mathcal{P}(\mathbb{Z}^{+})$, $I\neq J$ implies $L_{I}\neq L_{J}$.
@@ 0.4.2 Degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K4}_{t})$
Let $\mathbb{F}\in\mathsf{GF}_{r}(\mathsf{K4}_{t})$. Then (i) $\mathbb{F}\models\Diamond\top\vee\blacklozenge\top$ if and only if $\mathbb{F}\ncong\bullet$; (ii) $L^{*}=\mathsf{K}\oplus(\Diamond\top\vee\blacklozenge\top)$; and (iii) for all $L\in\mathsf{NExt}(\mathsf{K4}_{t})$, $L\subsetneq L^{*}$ if and only if $L=\mathsf{K4}_{t}$.
@@ 0.5 Kripke-incompleteness in $\mathsf{NExt}(\mathsf{S4}_{t})$
In this section, we focus on the degree of Kripke-incompleteness in $\mathsf{NExt}({\mathsf{S4}_{t}})$. In what follows, we write $\mathsf{deg}(L)$ for the degree of Kripke-incompleteness of $L$ in $\mathsf{NExt}({\mathsf{S4}_{t}})$. Our aim is to prove the dichotomy theorem for $\mathsf{NExt}({\mathsf{S4}_{t}})$. We first characterize the iterated splittings in $\mathsf{NExt}({\mathsf{S4}_{t}})$ and show that every iterated splitting has the degree of Kripke-incompleteness $1$. Then by constructing some frames generalizing the Nishimura-Rieger ladder, we show that all other tense logics in $\mathsf{NExt}({\mathsf{S4}_{t}})$ are of the degree of Kripke-incompleteness $2^{\aleph_{0}}$, which gives a characterization theorem for the degree of Kripke-incompleteness in $\mathsf{NExt}({\mathsf{S4}_{t}})$ and completes the proof of the dichotomy theorem.
@@
Recall from [13, Theorem 21] that ${\langle\mathsf{Log}(\mathfrak{Ch}_{2}),\mathsf{S5}_{t}\rangle}$ and ${\langle\mathsf{Log}(\mathfrak{Ch}_{1}),\mathscr{L}_{t}\rangle}$ are the only two splitting pairs in $\mathsf{NExt}({\mathsf{S4}_{t}})$. Then ${\{\mathscr{L}_{t},\mathsf{S5}_{t},{\mathsf{S4}_{t}}\}}$ is exactly the set of union-splittings in $\mathsf{NExt}({\mathsf{S4}_{t}})$ and so every union-splitting has the degree of Kripke-incompleteness $1$. However, since there are countably many tense logics in $\mathsf{NExt}(\mathsf{S5}_{t})$, Lemma 0.5.2 indicates that logics of the degree of Kripke-incompleteness $1$ are not necessary union-splittings, which shows that Blok’s characterization of the degree of Kripke-incompleteness for $\mathsf{NExt}(\mathsf{K})$ can not be generalized to $\mathsf{NExt}({\mathsf{S4}_{t}})$.
@@
For (1), recall first that $\mathfrak{F}_{L},w_{L}\not\models\varphi_{L}$. Then there exists a valuation $V^{\prime}:\mathsf{Prop}\to\mathcal{P}(X_{L})$ such that $\mathfrak{F}_{L},V^{\prime},w_{L}\models\neg\varphi_{L}$. Note that $\mathbb{F}_{I}$ is differentiated, there exists a valuation $V$ in $\mathbb{F}_{I}$ such that the following conditions hold: (i) $V{\upharpoonright}X_{L}=V^{\prime}$, (ii) $V(p)=R[b_{k}]\cup{\{x_{0},x_{1},x_{2},y_{0},y_{1}\}}$ and (iii) $V(q_{i})=R[b_{i}]$ for all $i\leq k$. Let $\mathfrak{M}=(\mathbb{F}_{I},V)$. Then $\mathfrak{M},b_{k+1}\models\varphi_{0}$ and $\mathfrak{M},w_{L}\models\neg\varphi_{L}$. Since $\mathbb{F}_{I}$ is $k$-transitive, $\mathfrak{M},x_{0}\models\nabla^{3}\mathsf{bd}_{k}$ and there exists a $p$-path from $x_{0}$ to $b_{k+1}$ of length $4$, we see that $\mathfrak{M},x_{0}\models\varphi_{x_{0}}$. Moreover, note that $\mathfrak{M},r_{0}\models\neg p$ and $V(\varphi_{0})\subseteq Y_{I}$, we have $\mathfrak{M},u\models\Delta^{3}\varphi_{0}\vee\neg\nabla_{p}^{4}\varphi_{0}$ for each point $u\in X_{I}\setminus{\{x_{0}\}}$. Thus $x_{0}$ is the unique point satisfying $\varphi_{x_{0}}$. By the construction the formulas $\varphi_{u}$, the reader can now easily check that $V(\varphi_{AB})=A\cup B$ and $\mathfrak{M},u\models\varphi_{u}$ for all $u\in U$. Since $w_{L}\in R_{\sharp}^{k}[u]$, we have $\mathfrak{M},u\models\varphi_{u}\wedge\Delta^{k}\neg\varphi_{L}$.
@@
For (4), take any $j\not\in I$. Suppose $\mathbb{F}_{I}\not\models\neg\varphi_{c_{j}}$. Then there exists $v\in X_{I}$ and a valuation $V$ in $\mathbb{F}_{I}$ such that $\mathbb{F}_{I},V,v\models\varphi_{c_{j}}$. Thus $V(\varphi_{x_{0}})\neq\varnothing$. By (2),$V(\varphi_{AB})=A\cup B$ and $V(\varphi_{w})={\{w\}}$ for $w\in{\{a_{j},a_{j+1}\}}$. Since $j\not\in I$, we have $\breve{R}[a_{j}]\setminus(A\cup B\cup\breve{R}[a_{j+1}])=\varnothing$ and so $\mathfrak{M}\models\neg\varphi_{c_{j}}$, which is a contradiction. ∎
@@
For all $I,J\in\mathcal{P}(\mathbb{Z}^{+})$, $I\neq J$ implies $L_{I}\neq L_{J}$.

### v2

@@ 0.4.1 Degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K}_{t})$
Since $\bullet\not\models\Diamond\top\vee\blacklozenge\top$ and $L^{*}=\mathsf{K}_{t}/\mathsf{Log}(\bullet)$, we get $L^{*}\subseteq\mathsf{K}\oplus(\Diamond\top\vee\blacklozenge\top)$. Take any $\mathbb{F}\in\mathsf{GF}_{r}(L^{*})$. Then $\mathbb{F}\ncong\bullet$. By Proposition 0.4.1, $\mathbb{F}\models\Diamond\top\vee\blacklozenge\top$. Hence $\Diamond\top\vee\blacklozenge\top\in\mathsf{Log}(\mathsf{GF}_{r}(L^{*}))=L^{*}$. ∎
@@
[…]
@@
For (1), we prove by induction on $n$. Let $n=0$. Note that $\breve{R_{I}}[0]=\varnothing$, $k\in R_{I}^{k}[0]$ and $\breve{R_{I}}^{k+1}[k]=\varnothing$, we have $\mathbb{F}_{I},0\models\gamma_{0}$. Suppose $\mathbb{F}_{I},x\models\gamma_{0}$. Then $\mathbb{F}_{I},x\models\Diamond^{k}\blacksquare^{k+1}\bot$, which entails that there exists a strict chain $\langle x_{i}:i\leq k\rangle$} with $x=x_{0}$. Since $|\mathfrak{F}_{L}|<k$, we see $x\not\in X_{L}$. Since $\mathbb{F}_{I},x\models\blacksquare\bot\wedge\Diamond\blacksquare^{2}\bot$, we have $x\not\in\mathbb{Z}^{+}\cup{\{i^{*}:i\in I\}}$. Thus $x=0$. Let $n>0$. By induction hypothesis, for all $y\in X_{I}$, $\mathbb{F}_{I},y\models\gamma_{n-1}$ if and only if y=n1y=n-1. Since $n-1\in\breve{R_{I}}[n]\setminus\breve{R_{I}}^{2}[n]$, we have $\mathbb{F}_{I},n\models\gamma_{n}$. Suppose $\mathbb{F}_{I},x\models\gamma_{n}$. By induction hypothesis, $x\in R_{I}[n-1]\setminus R_{I}^{2}[n-1]$, which entails $x=n$.
@@
Suppose there exists $\mathfrak{G}=(Y,S)\in\mathsf{Fr}(\mathsf{Log}(\mathbb{F}_{I}))$. By an easy induction, we see that for all $V\in A_{I}$, $V\cap\omega$ and $V\cap{\{i^{*}:i\in I\cup{\{0\}}\}}$ are either finite or co-finite. Thus $\mathbb{F}_{I},1\models\Box(\Box(p\to\Box p)\to p)\to\Box p$. By Lemma 0.4.10(1), $\mathbb{F}_{I}\models\gamma_{i}\to\Diamond\gamma_{i+1}$ for all $i\in\omega$. Thus we have
$\mathfrak{G}\models\gamma_{1}\to(\Box(\Box(p\to\Box p)\to p)\to\Box p)$ and $\mathfrak{G}\models{\{\gamma_{i}\to\Diamond\gamma_{i+1}:i\in\omega\}}$.
Note that $\mathbb{F}_{I}\models\Delta^{k}\gamma_{1}$ and $\gamma_{1}$ is variable-free, there exists $y\in Y$ such that $\mathfrak{G},y\models\gamma_{1}$. Since $\mathfrak{G}\models{\{\gamma_{i}\to\Diamond\gamma_{i+1}:i\in\omega\}}$, we see that there exists an infinite strictly ascending $S$-chain $\langle u_{i}:i\in\mathbb{Z}^{+}\rangle$} such that $y=u_{1}$ and $\mathfrak{G},u_{i}\models\gamma_{i}$ for all $i\in\mathbb{Z}^{+}$. Let $U$ be a valuation in $\mathfrak{G}$ such that $U(p)={\{u_{2i}:i\in\mathbb{Z}^{+}\}}$. Then we see that $\mathfrak{G},U,y\not\models\Box(\Box(p\to\Box p)\to p)\to\Box p$. Hence $\mathfrak{G}\not\models\gamma_{1}\to(\Box(\Box(p\to\Box p)\to p)\to\Box p)$, which contradicts $\mathfrak{G}\in\mathsf{Fr}(\mathsf{Log}(\mathbb{F}_{I}))$. ∎
@@
For all $\in\mathcal{P}(\mathbb{Z}^{+}), IJI\neq J$ implies $L_{I}\neq L_{J}$.
@@ 0.4.2 Degree of Kripke-incompleteness in $\mathsf{NExt}(\mathsf{K4}_{t})$
Let $\mathbb{F}\in\mathsf{GF}_{r}(\mathsf{K4}_{t})$. Then (i) $\mathbb{F}\models\Diamond\top\vee\blacklozenge\top$ if and only if $\mathbb{F}\ncong\bullet$; (ii) $L^{*}=\mathsf{K4}\oplus(\Diamond\top\vee\blacklozenge\top)$; and (iii) for all $\in\mathsf{NExt}(\mathsf{K4}_{t}), LL$$\subsetneq L^{*}$ if and only if $L=\mathsf{K4}_{t}$.
@@ 0.5 Kripke-incompleteness in $\mathsf{NExt}(\mathsf{S4}_{t})$
In this section, we focus on the degree of Kripke-incompleteness in $\mathsf{NExt}({\mathsf{S4}_{t}})$. In what follows, we write $\mathsf{deg}(L)$ for the degree of Kripke-incompleteness of $L$ in $\mathsf{NExt}({\mathsf{S4}_{t}})$. Our aim is to prove the dichotomy theorem for $\mathsf{NExt}({\mathsf{S4}_{t}})$. We first characterize the iterated splittings in $\mathsf{NExt}({\mathsf{S4}_{t}})$ and show that every iterated splitting has the degree of Kripke-incompleteness 11. Then by constructing some frames generalizing the Nishimura-Rieger ladder, we show that all other tense logics in $\mathsf{NExt}({\mathsf{S4}_{t}})$ are of the degree of Kripke-incompleteness $2^{\aleph_{0}}$, which gives a characterization theorem for the degree of Kripke-incompleteness in $\mathsf{NExt}({\mathsf{S4}_{t}})$ and completes the proof of the dichotomy theorem.
@@
Recall from [13, Theorem 21] that $\langle\mathsf{Log}(\mathfrak{Ch}_{2}),\mathsf{S5}_{t}\rangle$} and $\langle\mathsf{Log}(\mathfrak{Ch}_{1}),\mathscr{L}_{t}\rangle$} are the only two splitting pairs in $\mathsf{NExt}({\mathsf{S4}_{t}})$. Then $\{\mathscr{L}_{t},\mathsf{S5}_{t},{\mathsf{S4}_{t}}\}$} is exactly the set of union-splittings in $\mathsf{NExt}({\mathsf{S4}_{t}})$ and so every union-splitting has the degree of Kripke-incompleteness 11. However, since there are countably many tense logics in $\mathsf{NExt}(\mathsf{S5}_{t})$, Lemma 0.5.2 indicates that logics of the degree of Kripke-incompleteness 11 are not necessary union-splittings, which shows that Blok’s characterization of the degree of Kripke-incompleteness for $\mathsf{NExt}(\mathsf{K})$ can not be generalized to $\mathsf{NExt}({\mathsf{S4}_{t}})$.
@@
For (1), recall first that $\mathfrak{F}_{L},w_{L}\not\models\varphi_{L}$. Then there exists a valuation $V^{\prime}:\mathsf{Prop}\to\mathcal{P}(X_{L})$ such that $\mathfrak{F}_{L},V^{\prime},w_{L}\models\neg\varphi_{L}$. Note that $\mathbb{F}_{I}$ is differentiated, there exists a valuation $V$ in $\mathbb{F}_{I}$ such that the following conditions hold: (i) $V{\upharpoonright}X_{L}=V^{\prime}$, (ii) $V(p)=R[b_{k}]\cup{\{x_{0},x_{1},x_{2},y_{0},y_{1}\}}$ and (iii) $V(q_{i})=R[b_{i}]$ for all $i\leq k$. Let $\mathfrak{M}=(\mathbb{F}_{I},V)$. Then $\mathfrak{M},b_{k+1}\models\varphi_{0}$ and $\mathfrak{M},w_{L}\models\neg\varphi_{L}$. Since $\mathbb{F}_{I}$ is $k$-transitive, $\mathfrak{M},x_{0}\models\nabla^{3}\mathsf{bd}_{k}$ and there exists a $p$-path from $x_{0}$ to $b_{k+1}$ of length 44, we see that $\mathfrak{M},x_{0}\models\varphi_{x_{0}}$. Moreover, note that $\mathfrak{M},r_{0}\models\neg p$ and $V(\varphi_{0})\subseteq Y_{I}$, we have $\mathfrak{M},u\models\Delta^{3}\varphi_{0}\vee\neg\nabla_{p}^{4}\varphi_{0}$ for each point $u\in X_{I}\setminus{\{x_{0}\}}$. Thus $x_{0}$ is the unique point satisfying $\varphi_{x_{0}}$. By the construction the formulas $\varphi_{u}$, the reader can now easily check that $V(\varphi_{AB})=A\cup B$ and $\mathfrak{M},u\models\varphi_{u}$ for all $u\in U$. Since $w_{L}\in R_{\sharp}^{k}[u]$, we have $\mathfrak{M},u\models\varphi_{u}\wedge\Delta^{k}\neg\varphi_{L}$.
@@
[…]
@@
For all $\in\mathcal{P}(\mathbb{Z}^{+}), IJI\neq J$ implies $L_{I}\neq L_{J}$.